# Lattice Hamiltonians and Stray Interactions Within Quantum Processors

### v2

@@ IV $ZZ$ gate in a triangle
Contrary to common belief, which regards three-body interactions as weaker and often negligible compared to two-body interactions, this section emphasizes the need to consider three-body interactions in circuits. This consideration helps enhance two-qubit gate fidelity, furthermore it may pave the road toward introducing meaningful instantaneous three-qubit gates. It’s useful to categorize decoupling strategies into two distinct types: (1) strict isolation –this method aims to find an optimal operating point where the qubits involved in the gate operation are completely isolated, preventing even weak interactions with ungated qubits. The goal is to reach a state in which the coupling strength between the gated qubits and any additional, non-participating qubit is effectively null. (2) Soft isolation – Rather than striving for complete isolation, we may focus on reducing stray couplings.
A critical aspect of our research involves conducting an extensive analysis of three-body lattice interactions among the qubits and assessing their impact on the performance of two-qubit gates. Each strategy presents its own set of advantages and challenges. The choice between strict and soft decoupling often depends on the specific requirements of the quantum circuit and the nature of the qubits involved. We will understand in this section that the concept of strict isolation might not always be feasible or even advantageous. Thus, our aim will be to concentrate on mitigating the effects of these interactions instead of trying to eliminate them completely. This method is particularly useful in complex quantum circuits where perfect isolation is challenging to achieve, or where selective interactions can contribute positively to the system’s functionality.
Figure 5: Stray couplings in the triangular circuit with all-to-all connectivity as a function of the frequencies of the resonators C23 and C13 i.e. $\omega_{c2}$ and $\omega_{c3}$ respectively. The first row is at $g_{12}/2\pi=4$ MHz with nearly zero $J_{12}$ and $J_{13}$ at spot A with coupler frequencies $(\omega_{c12},\omega_{c23},\omega_{c13})/2\pi=(6.66,5.299,6.141)$ GHz, B: $(6.67,5.405,6.15)$ GHz and C: $(6.68,5.275,6.154)$ GHz. The second row is at $g_{12}/2\pi=13$ MHz with nearly zero $J_{12}$ and $J_{13}$ at spot D: $(5.438,5.317,6.021)$ GHz, E: $(5.443,5.472,6.076)$ GHz and F: $(5.446,5.621,6.098)$ GHz. The shaded regions denote two intersections: $|ZZI|\displaystyle\cap|ZIZ|<50$ kHz (light red) and ${\rm max}(|ZZI|,|ZIZ|,|ZZZ|)<50$ kHz (light green), both on top of the $IZZ$ interaction. Other circuit parameters are the same as in Fig. 4 with $\omega_{2}/2\pi=5.0$ GHz.
IV.1 Strict $J$-isolation
@@
The common belief is that the optimum operating point for a two-qubit gate within a triangular circuit is where two couplings are effectively turned off (‘OFF’), while the third remains active (‘ON’), thus establishing an ‘OFF-OFF-ON’ state. To analyze the operating point of this 2-qubit gate in presence of a third interacting qubit, first we begin by setting $J_{12}^{00}$ to zero, achievable at certain frequencies of the coupler $C_{12}$. Figure 5 showcases circuits with parameters outlined in the caption, all configured such that $J_{12}^{00}$ is nullified. We then focus on turning off the second coupling, say $J_{13}^{00}$, with its values indicated on the right axes of the figure. Simultaneously, the third interaction, $J_{23}^{00}$, marked on the upper axes, is selected to be strong. The variation of these couplings is achieved by adjusting the frequencies of couplers $C_{23}$ and $C_{13}$, as shown on the lower and left axes respectively.
@@
We evaluate the interaction strengths of the $ZZ$ and $ZZZ$ for the condition of $J_{12}^{00}=0$ , ensuring $Q1$ and $Q2$ do not interact. In Fig. 5 we achieve the isolation by testing two different values for direct capacitive coupling between $Q1$ and $Q2$, i.e. $g_{12}/{2\pi}=4,13$ MHz in the upper and lower rows, respectively. Given the circuit parameters selected, the next coupling parameter to be neutralized is $J_{13}^{00}=0$. The goal of this setting is to either preserve or enhance the coupling between $Q2$ and $Q3$, represented by $J_{23}^{00}$. This is achieved by tuning the frequency of the $C_{13}$ coupler within a range of 5.4 GHz to 7 GHz. As a result, the coupling strength $J_{13}/2\pi$ varies from $-3$ MHz to $+9$ MHz, reaching zero at the midpoint. Consequently, any point along the $J_{13}=0$ line is characterized by a precise OFF-OFF interaction, which is critical for the intended gate operation.
@@
In exploring the stray $ZZ$ and $ZZZ$ couplings, we focus on the horizontal line corresponding to $J_{13}=0$ in Fig. 5. This line signifies OFF-OFF interaction points achieved by setting both $J_{12}^{00}$ and $J_{13}^{00}$ to zero. While this condition is intended to decouple $Q1$ from $Q2$ and $Q3$, the figure illustrates that residual stray couplings still exist between $Q1$ and the other two qubits, notably in $ZZZ$ and $ZZI$ configurations. This observation is crucial as it demonstrates that strict isolation of seemingly non-interacting qubits does not necessarily lead to the elimination of stray couplings.
The persistence of these stray couplings is attributed to a significant factor. Although decoupling of two qubits, labeled $i$ and $j$, is achieved by setting their $J_{ij}^{00}$ (computational level coupling) to zero, these qubits remain indirectly coupled at their non-computational levels. These levels include couplings such as $J_{ij}^{01}$, $J_{ij}^{10}$, and others. It is these higher-order couplings that the stray interactions depend upon. Thus, even when qubits are seemingly decoupled at their fundamental computational states, indirect couplings at other energy levels persist, leading to stray interactions that must be considered in quantum circuit design and operation.
@@
The upper and lower plots in this column represent scenarios of weak and strong capacitive coupling between $Q1$ and $Q2$, denoted as $g_{12}$. In the case of weak $g_{12}$, all red dashed boundaries converge, forming an extensive frequency range conducive to effectively decoupling $Q1$ from the $Q2$-$Q3$ pair without significant stray couplings. This finding is particularly relevant for quantum devices designed for precise two-qubit gate operations, like the Controlled-$Z$ ($CZ$) gate. Devices with such characteristics can smoothly transition from low to high $ZZ$ interaction values while keeping spectator errors — errors affecting qubits not directly involved in the gate operation — to a minimum. This aspect is crucial for improving the overall performance and reliability of quantum gates in computational tasks.
@@
This observation reveals an important secret: assuming that three-body $ZZZ$ interactions are inherently weaker than two-body $ZZ$ interactions can lead to significant oversighting in gate design. In some scenarios, a circuit design might effectively suppress parasitic two-body $ZZ$ interactions but, inadvertently, introduce substantial error rates due to strong three-body interactions among qubits that are presumed to be noninteracting. Therefore, an accurate assessment of three-body interactions is essential for a comprehensive analysis of gate performance. Neglecting their potential impact could result in flawed assumptions and designs, ultimately hindering the achievement of high-fidelity quantum operations in certain quantum computing scenarios.
@@
Figure 6: Stray couplings in the triangular circuit with all-to-all connectivity as a function of the frequencies of the resonators C23 and C13 $\omega_{c23}$ and $\omega_{c13}$ respectively. The first row is at $g_{12}/2\pi=4$ MHz with weak $J_{12}/2\pi\approx 1$ MHz at $\omega_{c12}/2\pi=6.2$ GHz. The second row is at $g_{12}/2\pi=16$ MHz with weak $J_{12}/2\pi\approx 1$ MHz at $\omega_{c12}/2\pi=5.3$ GHz. Other circuit parameters are the same as Fig. 4 with $\omega_{2}/2\pi=5.0$ GHz.
IV.2 Soft $J$-isolation
@@
The common desire in multi-qubit circuit design is that during the application of a two-qubit gate, other qubits are silenced. This is usually interpreted as turning off $J$ interaction between the gated qubits and other qubits. In the previous section we examine this concept in the context of a triangular lattice by setting two pairwise interactions to zero and amplifying the third. We referred to this concept as ‘strict $J$-isolation’. Our more detailed analysis reveals that the situation is significantly more complex than nullifying $J$ interaction.
@@
In this section, we explore the same OFF-OFF-ON interaction between $Q2$ and $Q3$ in an alternative setup: Rather than enforcing a complete absence of interaction of these two qubits with $Q1$, we allow some weak $J$-interactions to take place if they contribute to reducing parasitic interactions. Our goal shifts towards minimizing stray coupling strengths to the extent that they either disappear or become negligible (i.e., less than 50 kHz). Enforcing zero interaction between qubits is limited in scope; it only prevents interactions between two specific energy levels of one qubit with another, which is inadequate to prohibit transitions across all energy levels. For instance, while strict decoupling might set $J_{12}^{00}$ to zero, other coupling strengths like $J_{12}^{01}$ could still be non-zero.
@@
This realization implies that the idea of completely nullifying $J$-interaction between qubits might be overly simplistic, if not altogether unachievable. Recognizing this limitation can shift the focus towards managing stray interactions rather than striving for complete non-interaction. In other words, we focus on the impact of relaxing one of the two OFF couplings to a near-OFF state, denoted as $\widetilde{\textup{OFF}}$. This softer approach to qubit isolation, or ‘soft isolation’, is explored as a potentially more effective and realistic strategy for enhancing $ZZ$ interaction between $Q2$ and $Q3$, as opposed to the strict decoupling strategy previously considered.
@@
Soft Isolation Criterion: By setting the interaction between $Q1$ and $Q2$ to $J_{12}^{00}/2\pi=1$ MHz, rather than completely zeroing it, we introduce a ‘soft isolation’ criterion. Figure 6 illustrates how this soft isolation impacts both two-body and three-body stray couplings in two different circuit configurations: First (Second) row denotes a circuit with a weak (strong) direct capacitive coupling between $Q1$ and $Q2$, $g_{12}=4(16)$ MHz. Such strenth occurs at $\omega_{C_{12}}/2\pi=6.2(5.3)$ GHz.
@@
The identification of ‘safe zones’ for low stray couplings is crucial, which is identifiable in the last column of Fig. 6. These zones are defined by regions where all stray interactions are under 50 kHz."The region marked within the dashed green boundary in the final column presents zones exhibiting minimal stray couplings, both two-body and three-body. Upon comparison with scenarios of strict isolation, it becomes evident that in the example of larger direct qubit-qubit coupling $g_{12}$, the zone of minimal stray interaction—termed the ‘stray safe zone’—extends over a broader area in conditions of soft isolation than in those of strict isolation. This suggests that enhances the practical applicability of soft isolation techniques in managing stray couplings.
@@
A comparative analysis with strict decoupling shows that soft decoupling scenario can offer equivalent or broader stray-free zones compared to the strict decoupling scenario. This is a highly non-trivial and significant insight that our lattice gate theory is able to predict. In brief, a more flexible approach to qubit isolation can lead to better control and optimization of quantum gate operations. These findings are valuable for advancing quantum computing technology, particularly in optimizing multi-qubit circuits for high-fidelity operations.
@@ V $ZZZ$ Superiority
We study the two-body and three-body stray interactions in superconducting quantum processors, with a particular focus on scenarios where the strength of three-body interactions $ZZZ$ surpasses that of two-body interactions. This observation has significant implications for quantum computing, particularly in the design of quantum gates.
@@
Figure 7: Absolute ratio of the three-body $ZZZ$ and the maximum of the two-body interactions occurring in the triangular circuit in Fig. 3 as a function of the frequencies of resonators C23 and C13 $\omega_{c23}$ and $\omega_{c13}$ at (a) $g_{12}/2\pi=4$ MHz and (b) $g_{12}/2\pi=13$ MHz, respectively. The green contour line represents $|\alpha_{ZZZ}|={\rm max}(|\alpha_{Z_{i}Z_{j}I_{k}}|)$. (c) Stray couplings as a function of C3 frequency at $\omega_{c23}/2\pi=5.272$ GHz (dashed white line on the top plots). Gray area indicate stray couplings are less than 50 kHz, light green area show the validity of $ZZZ$ superiority. Other circuit parameters are the same as in Fig. 4 with $\omega_{2}/2\pi=5.0$ GHz.
@@
Circles labeled A to F in Fig. 7(a,b) represent three sweet spots for the OFF-OFF-ON gate studied in Fig. 5. The analysis indicates that in certain ON spots which are near the three-body superiority zone, such as point A, are not ideal as a safe gate sweet spot. In such operational points, while $ZZ$ strengths are small, the three-body $ZZZ$ interaction can exceed $ZZ$’s, which can lead to noisy gate operations. Figure 7 (c, d) show the behavior of stray couplings at the fixed frequency $\omega_{C_{23}}/2\pi=5.272$ GHz, represented by a white dashed line in Fig. 7 (a, b). The behavior is examined across a range of $\omega_{C_{23}}$ frequencies for both weak and strong $g_{12}$. In the weak direct coupling $g_{12}$ case, a narrow band around $\omega_{c13}/2\pi=6$ GHz can be identified in which all residual stray interactions are small. This observation indicates that the specified frequency range may act as a zone with minimal error, optimal for establishing a quantum idle point. Consequently, this parameter domain is suitable for implementing a Controlled-$Z$ ($CZ$) gate between qubits Q2 and Q3.

### v3

@@ IV $ZZ$ gate in a triangle
Contrary to common belief, which regards three-body interactions as weaker and often negligible compared to two-body interactions, this section emphasizes the need to consider three-body interactions in circuits. This consideration helps enhance two-qubit gate fidelity, furthermore it may pave the road toward introducing meaningful instantaneous three-qubit gates. It’s useful to categorize decoupling strategies into two distinct types: (1) hard decoupling –this method aims to find an optimal operating point where the qubits involved in the gate operation are completely isolated, preventing even weak interactions with ungated qubits. The goal is to reach a state in which the coupling strength between the gated qubits and any additional, non-participating qubit is effectively null. (2) Soft decoupling – Rather than striving for strict decoupling, we may focus on reducing stray couplings.
A critical aspect of our research involves conducting an extensive analysis of three-body lattice interactions among the qubits and assessing their impact on the performance of two-qubit gates. Each strategy presents its own set of advantages and challenges. The choice between hard and soft decoupling often depends on the specific requirements of the quantum circuit and the nature of the qubits involved. We will understand in this section that the concept of hard decoupling might not always be feasible or even advantageous. Thus, our aim will be to concentrate on mitigating the effects of these interactions instead of trying to eliminate them completely. This method is particularly useful in complex quantum circuits where strict decoupling is challenging to achieve, or where selective interactions can contribute positively to the system’s functionality.
IV.1 Hard $J$-decoupling
@@
(a)(b)
Figure 5: (a) Hard $J$-decoupling at $J_{12}=J_{13}\approx 0$, (b) Soft $J$-decoupling at $J_{12}\sim$ 1 MHz and $J_{13}=0$
The common belief is that the optimum operating point for a two-qubit gate within a triangular circuit is where two couplings are effectively turned off (‘OFF’), while the third remains active (‘ON’), thus establishing an ‘OFF-OFF-ON’ state. To analyze the operating point of this 2-qubit gate in presence of a third interacting qubit, first we begin by setting $J_{12}^{00}$ to zero, achievable at certain frequencies of the coupler $C_{12}$. This is what is namely ‘hard $J$-decoupling’, which is diagrammatically sketched in Fig. 5(a). In our simulation we apply the condition step by step, first we turn off coupling between Q1 and Q2, then between Q1 and Q3, and finally switch on and off $CZ$ gate between Q2 and Q3. Figure 6 showcases circuits with parameters outlined in the caption, all configured such that $J_{12}^{00}$ is nullified. We then focus on turning off the second coupling, say $J_{13}^{00}$, with its values indicated on the right axes of the figure. Simultaneously, the third interaction, $J_{23}^{00}$, marked on the upper axes, is selected to be strong. The variation of these couplings is achieved by adjusting the frequencies of couplers $C_{23}$ and $C_{13}$, as shown on the lower and left axes respectively.
Figure 6: Stray couplings in the triangular circuit with all-to-all connectivity calculated by numerically diagonalizing the Hamiltonian Eq. (III), as a function of the frequencies of the resonators C23 and C13 i.e. $\omega_{c2}$ and $\omega_{c3}$ respectively. The first row is at $g_{12}/2\pi=4$ MHz with nearly zero $J_{12}$ and $J_{13}$ at spot A with coupler frequencies $(\omega_{c12},\omega_{c23},\omega_{c13})/2\pi=(6.66,5.299,6.141)$ GHz, B: $(6.67,5.405,6.15)$ GHz and C: $(6.68,5.275,6.154)$ GHz. The second row is at $g_{12}/2\pi=13$ MHz with nearly zero $J_{12}$ and $J_{13}$ at spot D: $(5.438,5.317,6.021)$ GHz, E: $(5.443,5.472,6.076)$ GHz and F: $(5.446,5.621,6.098)$ GHz. The shaded regions denote two intersections: $|ZZI|\displaystyle\cap|ZIZ|<50$ kHz (light red) and ${\rm max}(|ZZI|,|ZIZ|,|ZZZ|)<50$ kHz (light green), both on top of the $IZZ$ interaction. Other circuit parameters are the same as in Fig. 4 with $\omega_{2}/2\pi=5.0$ GHz.
@@
We evaluate the interaction strengths of the $ZZ$ and $ZZZ$ for the condition of $J_{12}^{00}=0$ , ensuring $Q1$ and $Q2$ do not interact. In Fig. 6 we achieve the decoupling by testing two different values for direct capacitive coupling between $Q1$ and $Q2$, i.e. $g_{12}/{2\pi}=4,13$ MHz in the upper and lower rows, respectively. Given the circuit parameters selected, the next coupling parameter to be neutralized is $J_{13}^{00}=0$. The goal of this setting is to either preserve or enhance the coupling between $Q2$ and $Q3$, represented by $J_{23}^{00}$. This is achieved by tuning the frequency of the $C_{13}$ coupler within a range of 5.4 GHz to 7 GHz. As a result, the coupling strength $J_{13}/2\pi$ varies from $-3$ MHz to $+9$ MHz, reaching zero at the midpoint. Consequently, any point along the $J_{13}=0$ line is characterized by a precise OFF-OFF interaction, which is critical for the intended gate operation.
@@
In exploring the stray $ZZ$ and $ZZZ$ couplings, we focus on the horizontal line corresponding to $J_{13}=0$ in Fig. 6. This line signifies OFF-OFF interaction points achieved by setting both $J_{12}^{00}$ and $J_{13}^{00}$ to zero. While this condition is intended to decouple $Q1$ from $Q2$ and $Q3$, the figure illustrates that residual stray couplings still exist between $Q1$ and the other two qubits, notably in $ZZZ$ and $ZZI$ configurations. This observation is crucial as it demonstrates that hard decoupling of seemingly non-interacting qubits does not necessarily lead to the elimination of stray couplings.
The persistence of these stray couplings is attributed to a significant factor. Although decoupling of two qubits, labeled $i$ and $j$, is achieved by setting their $J_{ij}^{00}$ (computational level coupling) to zero, these qubits remain indirectly coupled at their non-computational levels. These levels include couplings such as $J_{ij}^{01}$, $J_{ij}^{10}$, and others. It is these higher-order couplings that the stray interactions depend upon. Thus, even when qubits are seemingly decoupled at their fundamental computational states, indirect couplings at other energy levels persist, leading to stray interactions that must be considered in quantum circuit design and operation. Details can be found in Appendix E.
@@
The upper and lower plots in this column represent scenarios of weak and strong capacitive coupling between $Q1$ and $Q2$, denoted as $g_{12}$. In the case of weak $g_{12}$, all red dashed boundaries converge, forming an extensive frequency range conducive to effectively decoupling $Q1$ from the $Q2$-$Q3$ pair without significant stray couplings. This finding is particularly relevant for quantum devices designed for precise two-qubit gate operations, like the $CZ$ gate. Devices with such characteristics can smoothly transition from low to high $ZZ$ interaction values while keeping spectator errors — errors affecting qubits not directly involved in the gate operation — to a minimum. This aspect is crucial for improving the overall performance and reliability of quantum gates in computational tasks.
@@
This observation reveals that assuming three-body $ZZZ$ interactions are weaker than two-body $ZZ$ interactions can lead to significant oversights in gate design. Suppressing parasitic two-body $ZZ$ interactions may inadvertently introduce substantial errors from strong three-body interactions among presumed noninteracting qubits. Therefore, accurately assessing three-body interactions is essential to avoid flawed designs and achieve high-fidelity quantum operations.
@@
Figure 7: Stray couplings in the triangular circuit with all-to-all connectivity calculated by numerically diagonalizing the Hamiltonian Eq. (III), as a function of the frequencies of the resonators C23 and C13 $\omega_{c23}$ and $\omega_{c13}$ respectively. The first row is at $g_{12}/2\pi=4$ MHz with weak $J_{12}/2\pi\approx 1$ MHz at $\omega_{c12}/2\pi=6.2$ GHz. The second row is at $g_{12}/2\pi=16$ MHz with weak $J_{12}/2\pi\approx 1$ MHz at $\omega_{c12}/2\pi=5.3$ GHz. Other circuit parameters are the same as Fig. 6.
IV.2 Soft $J$-decoupling
@@
The common desire in multi-qubit circuit design is that during the application of a two-qubit gate, other qubits are silenced. This is usually interpreted as turning off $J$ interaction between the gated qubits and other qubits. In the previous section we examine this concept in the context of a triangular lattice by setting two pairwise interactions to zero and amplifying the third. In the previous section we referred to this concept as hard $J$-decoupling. However, our detailed analysis reveals that the situation is significantly more complex than nullifying $J$ interaction.
@@
In this section, we explore the same OFF-OFF-ON interaction between $Q2$ and $Q3$ in an alternative setup: Rather than enforcing a complete absence of interaction of these two qubits with $Q1$, we allow some weak $J$-interactions to take place if they contribute to reducing parasitic interactions. Our goal shifts towards minimizing stray coupling strengths to the extent that they either disappear or become negligible (i.e., less than 50 kHz). Enforcing zero interaction between qubits is limited in scope; it only prevents interactions between two specific energy levels of one qubit with another, which is inadequate to prohibit transitions across all energy levels. For instance, while hard decoupling might set $J_{12}^{00}$ to zero, other coupling strengths like $J_{12}^{01}$ could still be non-zero. The so-called ‘soft $J$-decoupling’, is sketched diagrammatically in Fig. 5(b).
@@
This realization implies that the idea of completely nullifying $J$-interaction between qubits might be overly simplistic, if not altogether unachievable. Recognizing this limitation can shift the focus towards managing stray interactions rather than striving for complete non-interaction. In other words, we focus on the impact of relaxing one of the two OFF couplings to a near-OFF state, denoted as $\widetilde{\textup{OFF}}$. This softer approach to qubit decoupling, or ‘soft $J$-decoupling’, is explored as a potentially more effective and realistic strategy for enhancing $ZZ$ interaction between $Q2$ and $Q3$, as opposed to the hard decoupling strategy previously considered.
@@
Soft $J$-decoupling Criterion: By setting the interaction between $Q1$ and $Q2$ to $J_{12}^{00}/2\pi=1$ MHz, rather than completely zeroing it, we introduce a ‘soft decoupling ’ criterion. Figure 7 illustrates how this can impact both two and three-body stray couplings in different circuit configurations: First (Second) row denotes a circuit with a weak (strong) direct capacitive coupling between $Q1$ and $Q2$, $g_{12}=4(16)$ MHz. Such strength occurs at $\omega_{C_{12}}/2\pi=6.2(5.3)$ GHz.
@@
The identification of ‘safe zones’ for low stray couplings is crucial, which is identifiable in the last column of Fig. 7. These zones are defined by regions where all stray interactions are under 50 kHz."The region marked within the dashed green boundary in the final column presents zones exhibiting minimal stray couplings, both two-body and three-body. Upon comparison with scenarios of hard decoupling, it becomes evident that in the example of larger direct qubit-qubit coupling $g_{12}$, the zone of minimal stray interaction—termed the ‘stray safe zone’—extends over a broader area in conditions of soft decoupling than in those of hard decoupling. This suggests that enhances the practical applicability of soft decoupling techniques in managing stray couplings.
@@
Refer to caption(b)
Figure 8: (a) $CZ$ gate fidelity between Q2 and Q3 for the cases of hard and soft decoupling, as a function of $\omega_{c23}$ and time $t$. (b) Maximum $CZ$ gate fidelity as a function of $\omega_{c23}$ for the cases of hard decoupling (blue) and soft decoupling (red).
To quantify the impact of three-body $ZZZ$ interactions on two-qubit gates, we evaluate the fidelity of the $CZ$ gate between Q2 and Q3 for both hard $J$ and soft $J$ decoupling scenarios. For simplicity, we assume no decoherence from the qubits to the environment and calculate the maximum fidelity as a function of the coupler frequencies $\omega_{c23}$ at specific coupler frequency $\omega_{c13}$ within the low-error zone. The comparison in Fig. 8 indicates that the $ZZZ$ interaction introduces approximately a 0.2% error, thus excluding the hard decoupling case beyond the error correction threshold.
Overall, our comparative analysis with hard decoupling demonstrates that the soft decoupling scenario can provide equivalent or broader stray-free zones compared to the hard decoupling scenario. The comparison indicates that higher-level interactions between gated and ungated qubits play a crucial role and require careful management. These characteristics can be encapsulated in the form of $ZZZ$ interactions. In quantum processors, where the fundamental operations are single- and two-qubit gates, these interactions should be carefully suppressed either at the design stage by reducing connectivity or lowering capacitive coupling, or at the gate stage by employing the soft decoupling method described above, which involves tuning circuit parameters such as the coupler frequency to optimize performance and suppress unwanted interactions. By integrating these strategies, we can enhance the reliability and efficiency of quantum processors.
@@ V $ZZZ$ Superiority
We study the two-body and three-body stray interactions in superconducting quantum processors, with a particular focus on scenarios where the strength of three-body interactions $ZZZ$ surpasses that of two-body interactions. This observation would potentially affect the design consideration of future superconducting transmon architectures.
@@
Figure 9: Absolute ratio of the three-body $ZZZ$ and the maximum of the two-body interactions occurring in the triangular circuit in Fig. 3 as a function of the frequencies of resonators C23 and C13 $\omega_{c23}$ and $\omega_{c13}$ at (a) $g_{12}/2\pi=4$ MHz and (b) $g_{12}/2\pi=13$ MHz, respectively. The green contour line represents $|\alpha_{ZZZ}|={\rm max}(|\alpha_{Z_{i}Z_{j}I_{k}}|)$. Corresponding stray couplings at specific coupler frequency $\omega_{c23}/2\pi=5.272$ GHz (dashed white line on the top plots) are shown in (c) and (d), respectively.
Gray area indicate stray couplings are less than 50 kHz, light green area show the validity of $ZZZ$ superiority. Other circuit parameters are the same as in Fig. 4 with $\omega_{2}/2\pi=5.0$ GHz.
@@
Circles labeled A to F in Fig. 9(a,b) represent three sweet spots for the OFF-OFF-ON gate studied in Fig. 6. The analysis indicates that in certain ON spots which are near the three-body superiority zone, such as point A, are not ideal as a safe gate sweet spot. In such operational points, while $ZZ$ strengths are small, the three-body $ZZZ$ interaction can exceed $ZZ$’s, which can lead to noisy gate operations. Figure 9 (c, d) show the behavior of stray couplings at the fixed frequency $\omega_{C_{23}}/2\pi=5.272$ GHz, represented by a white dashed line in Fig. 9 (a, b). The behavior is examined across a range of $\omega_{C_{23}}$ frequencies for both weak and strong $g_{12}$. In the weak direct coupling $g_{12}$ case, a narrow band around $\omega_{c13}/2\pi=6$ GHz can be identified in which all residual stray interactions are small. This observation indicates that the specified frequency range may act as a zone with minimal error, optimal for establishing a quantum idle point. Consequently, this parameter domain is suitable for implementing a Controlled-$Z$ ($CZ$) gate between qubits Q2 and Q3. Another example of a three-qubit circuit, where three qubits are coupled to a shared coupler, is provided in Appendix F.